Find the force between two point charges 5uc and 10uc separated by a distance of 0.2m in air


1
Expert's answer
2021-03-18T14:04:09-0400

The force between two point charges is Coulomb's force:


"F = k\dfrac{|q_1||q_2|}{r^2}"

where "q_1 = 5\mu C = 5\times 10^{-6}C, \space q_2 = 10\mu C = 10\times 10^{-6}C" are point charges, "k = 9\times 10^9 N\cdot m^2/C^2" is the Coulomb's constant, and "r= 0.2m" is the distance between the charges. Substituting these numbers, obtain:


"F = 9\times 10^{9}\dfrac{|5\times 10^{-6}||10\times 10^{-6}|}{0.2^2} =11.25N"

Answer. 11.25 N.